I am really confused, how do you multiply and divide fractions?
For example what is 6 divided by 2/3? and 2/5 x 1/2?

Answer by elima(1433)

Multiplying and dividing fractions are much easier than adding and subtracting.
6 divided by 2/3?
With your first problem, when it comes to dividing fractions, you need to multiply the inverseinvert the divisor:
Which is 2/3 and it becomes 3/2
Now we multiply the fractions;
6/1*3/2=18/2=9
2/5 x 1/2?
You multiply the numerator times the numerator and denominator times the denominator.
2/5*1/2=2*1, 5*2
2/10=1/5
Hope you understand
=)